An angle is 20 degree more than three times the given angle. If the two angles are supplementary, the angle are?
A) 20, 160 B) 40,140 C) 60,120 D) 70, 110

Answer by Cromlix(4381)

Given angle = x
Second angle = 20 + 3x
x + (20 + 3x) = 180
4x + 20 = 180
4x = 160
x = 40
So, given angle = 40 degrees.
Second angle = 140 degrees.
Answer 'B'
hope this helps.
:-)